I'm not sure how to solve this problem because I'm having trouble finding the equation. The question is - The points (3,4), (11,6), and (23,9) lie on the same line. Calculate the coordinates of another point on this line whose x coordinate is greater than 800.
Please help.

Answer by checkley71(8403)

(3,4) & (11,6) RESULT IN A SLOPE OF (6-4)/11-3) OR 2/8 OR .25 FOR THE SLOPE WHEN X=920 THEN Y=90
PROOF (Y-4)/(823-3)=.25 OR Y-4=800*.25 OR Y-4=205 OR Y=209 THUS THE DESIRED COORDINATE IS (823,209)